OpenStudy (anonymous):

Evaluate 3(a – 4b) when a = –2 and b = 5. A. 54 B. –54 C. 39 D. –66

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Replace a with -2 and b with 5 in the expression. Then do the arithmetic. 3(a – 4b) = = 3(-2 - 4 * 5)

OpenStudy (mathstudent55):

Start inside the parentheses and follow the correct order of operations.

OpenStudy (anonymous):

i'm not getting the answer

OpenStudy (anonymous):

can you help me @UsukiDoll

OpenStudy (anonymous):

-66 is the answer!!!!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Haven't you get it.... I worked it out in my head it's very easy....

OpenStudy (anonymous):

You are welcome!!

OpenStudy (anonymous):

And thank you for giving me a medal!!

OpenStudy (usukidoll):

" 3(a – 4b) = = 3(-2 - 4 * 5)" =3(-2-20) = 3(-22) = -66 yup that user is right.